# Light of Normals: Unified Feature Representation for Universal Photometric Stereo

### v2

@@ 5 Conclusion
A1.2.1 Categorization Methodology
To rigorously evaluate and enhance the capability of our LiNO-UniPS method for reconstructing surface normals of objects that feature high-frequency geometric details on complex surfaces, we curated a dedicated set of objects exhibiting diverse geometric complexities. These objects were subsequently graded by difficulty into five distinct levels, designated Level 1 to Level 5. Specifically, Levels 1–4 are classified following Dora [7] criterion, based on the number of salient edges $N_{\Gamma}$. Level 5, in contrast, is distinguished by the use of normal maps in its rendering. The specific criteria for this classification are as follows:
@@
A1.2.3 Lighting Setup
When rendering PS-Verse, we use four types of light sources; (a) environment lighting, (b) directional lighting, (c) point lighting, (d) uniform background lightings.
During rendering, we generate ten distinct lighting configurations by combining several base lighting components (conceptually denoted here as (a), (b), (c), and (d)). These specific configurations are as follows: (1) Component (a), (2) Component (b), (3) Component (b), (4) Components (a) + (b), (5) Components (a) + (c), (6) Components (b) + (c), (7) Components (a) + (b) + (c), (8) Components (a) + (d), (9) Components (b) + (d), (10) Components (a) + (b) + (d). The lighting setup includes: one directional light; a variable number of point lights, randomly ranging from one to three; and uniform background lighting, introduced to better simulate realistic global illumination. Every scene in PS-Verse is rendered as 20 images, each employing a lighting setup randomly chosen from our ten predefined lighting configurations. An example of such an image set for a single scene is illustrated in Figure A3.
A1.2.4 Comparison with Other Datasets
Here, we primarily compare several training datasets: CyclesPS-Train [25], PS-Wild [28], PS-Mix [31], PS-Uni MS-PS [20], and our PS-Verse. For a qualitative comparison of these datasets, please refer to Tab. 1. We now present illustrative qualitative comparisons in Fig. A4. Our comprehensive evaluation, encompassing both qualitative and quantitative aspects, leads us to conclude that PS-Verse is the premier training dataset in terms of quality for the Universal PS task.
@@
A1.3 Additional Discussion
A1.3.1 Why Feature Consistency Matters
@@
Building upon approaches such as UniPS [28] and SDM-UniPS [31] that utilize comparable decoder designs, a primary driver for our LiNO-UniPS is the introduction of an improved encoder. The objective of this encoder is to yield more consistent features, thereby achieving superior results. However, it is not an oversimplification that greater feature consistency directly translates to superior normal reconstruction; the decoder’s architecture and capabilities also represent a non-negligible factor in overall performance. Nonetheless, the decoders utilized in UniPS, SDM-UniPS and our LiNO-UniPS are architecturally similar, all adhering to the pixel-sampling paradigm. Consequently, when analyzing the relationship between feature consistency (e.g., as measured by CSIM/SSIM) and normal reconstruction performance in Fig. 1, we group these three methods together to facilitate a more direct comparison of the impact of their respective encoder-derived features.
@@
Ours w/MLP refers to a configuration of our LiNO-UniPS where the standard decoder is replaced by a simple two-layer MLP. As illustrated in Fig. A8, this variant gives the highest feature similarity. Visual inspection of the PCA plots further reveals that its extracted features have effectively disentangled lighting information. We hypothesize that the superior feature similarity of Ours w/mlp compared to the full LiNO-UniPS (with its original decoder) stems from the constraints imposed by the weaker MLP decoder; this limited decoder capacity compels the encoder to learn more consistent features to facilitate accurate normal reconstruction. While this enhanced feature consistency from the encoder may not entirely compensate for the reduced representational power of the simpler decoder in terms of absolute normal reconstruction quality (when compared to the full LiNO-UniPS), Ours w/MLP variant nevertheless significantly outperforms SDM-UniPS. This finding strongly corroborates our central hypothesis regarding the critical role of a powerful and well-regularized encoder in achieving effective feature disentanglement and consistency.
@@
Uni MS-PS also demonstrates high feature similarity. However, visual analysis of its features (Fig. A8) suggests that they remain considerably entangled with lighting information. Consequently, we infer that its high reported feature similarity may be more attributable to geometric self-consistency within its representations rather than successful illumination decoupling. Despite this apparent lack of complete feature decoupling, Uni MS-PS often achieves commendable reconstruction results. We attribute this primarily to its multi-scale architecture: beyond the initial stage, each subsequent network stage in Uni MS-PS incorporates predicted normals from the preceding stage as an additional input, effectively leveraging them as a strong geometric prior. This iterative refinement, guided by intermediate normal predictions, places Uni MS-PS in a distinct operational paradigm compared to methods like UniPS, SDM-UniPS, and our LiNO-UniPS.
@@
Furthermore, We need to figure out that Uni MS-PS exhibits certain practical limitations. (a) Its multi-scale nature leads to considerable inference latency, particularly when processing multiple high-resolution input images (e.g., handling 16 images at 4K resolution can extend to several hours). In contrast, our LiNO-UniPS method typically completes inference within tens of seconds for similar inputs. (b) While Uni MS-PS can reconstruct detailed surface normals, its reliance on potentially lower-resolution training datasets and its patch-based inference mechanism can lead to a loss of global contextual information, sometimes resulting in reconstructions that are locally detailed but globally inconsistent or erroneous 1.
A1.3.2 Light Registers Tokens
Figure A5: Attention maps for light registers tokens, showcasing their ability to focus on illumination-related regions within the input.
To achieve a more effective decoupling of lighting from normal features within our encoder, we introduce light register tokens. These tokens are specifically designed with the expectation that they will primarily capture and encode illumination information. Fig. A5 presents attention maps derived from the final layer of our encoder. These visualizations demonstrate that the light register tokens indeed attend to diverse illumination-related regions within the input images, including areas characterized by highlights and shading.
We introduce three additional light registration tokens: $x_{\text{hdri}}$ intended for environment light, $x_{\text{point}}$ for point light sources, and $x_{\text{area}}$ for area light sources. Fig. A6 shows the attention map visualizations for these different tokens. It is observable that the HDRI register typically attends to broader regions.
The point light register token and the area light register token exhibit similar attentional patterns, both demonstrating a more localized and sharper focus. This shared characteristic aligns well with the generally low-frequency nature of global HDRI illumination and the more localized, often high-frequency, impact typically associated with point and area light sources.
Figure A6: Attention maps for our different light registers tokens reveals distinct focusing patterns. The HDRI register token typically attends to broader, more diffuse regions, while the point light register token and the area light register token exhibit similar behaviors, both demonstrating a more localized and sharper attentional focus.
A1.3.3 Limitations
@@
Figure A7: For near-planar objects possessing intricate concave and convex surface details, our LiNO-UniPS tends to invert the predicted surface normals. The objects are from DiLiGenT-$\Pi$ [52]
@@
Despite the commendable performance demonstrated by LiNO-UniPS, certain limitations nonetheless remain, offering avenues for future research.
@@
Secondly, a limitation arises in the estimation of surface normals for near-planar objects that exhibit intricate concave and convex surface details, where the performance of LiNO-UniPS can be suboptimal. We attribute this primarily to a fundamental challenge inherent in the Universal Photometric Stereo (PS) paradigm: the absence of explicit light source parameters makes it difficult for the network to unambiguously determine the precise illumination direction (e.g., to distinguish between light originating from above versus below the surface). Consequently, this ambiguity can frequently, and in some instances, even lead to the estimated surface normals being inverted A7. In future investigations, we plan to explore whether more sophisticated lighting alignment strategies could effectively mitigate this specific issue.
A1.3.4 Broader Impacts
Our research offers several positive societal implications. It promises to drive technological advancements in digital content creation, including fields such as virtual reality (VR) and cinematic visual effects (VFX). Moreover, enhanced 3D reconstruction capabilities can bolster machine perception for applications like robotics and autonomous driving, and provide valuable tools for auxiliary scientific research.
Nevertheless, the advanced capabilities described also entail potential negative societal effects. High-fidelity 3D reconstruction, if misused, could be exploited to generate disinformation or fabricate convincing false realities. Furthermore, such technology raises concerns regarding potential infringements on personal privacy, and could pose risks to intellectual property rights and security.
A1.4 Extended Results
In this section, we present additional results to further demonstrate the capabilities of our LiNO-UniPS.
A1.4.1 Real Data
Fig. A9 shows reconstruction results on challenging real-world data, which was captured from diverse scenes, is mask-free, and features a 4K spatial resolution. These results demonstrate that our LiNO-UniPS is also scale, mask-free and detailed.
Figure A8: Post-PCA visualization of features extracted by different method encoders for the CowPNG from the DiLiGenT [48]. Metrics displayed to the right of each row is (CSIM/SSIM), higher values indicate higher feature similarity.
Figure A9: Top row: Example from the input multi-light images. Bottom row: Surface normal map reconstructed by our LiNO-UniPS. The data utilized is from SDM-UniPS [31], and the resolution is 4K .
Figure A10: Real-world data with masks and corresponding LiNO-UniPS reconstruction results; data sourced from UniPS [28] and SDM-UniPS [31].
@@
Fig. A11 presents a comparison of various Universal PS methods on our PS-Verse Testdata. Given that PS-Verse Testdata is a synthetic dataset, ground truth is readily available, facilitating precise quantitative evaluation. The results clearly demonstrate that our LINO-UniPS significantly outperforms contemporary approaches, including Uni MS-PS [20], SDM-UniPS [31], and UniPS [28].

### v4

@@ 5 Conclusion
A1.3.1 Categorization Methodology
To rigorously evaluate and enhance the capability of our LINO UniPS method for reconstructing surface normals of objects that feature high-frequency geometric details on complex surfaces, we curated a dedicated set of objects exhibiting diverse geometric complexities. These objects were subsequently graded by difficulty into five distinct levels, designated Level 1 to Level 5. Specifically, Levels 1–4 are classified following Dora Chen et al. (2024) criterion, based on the number of salient edges $N_{\Gamma}$. Level 5, in contrast, is distinguished by the use of normal maps in its rendering. The specific criteria for this classification are as follows:
@@
A1.3.3 Lighting Setup
When rendering PS-Verse, we use four types of light sources; (a) environment lighting, (b) directional lighting, (c) point lighting, (d) uniform background lighting.
During rendering, we generate ten distinct lighting configurations by combining several base lighting components (conceptually denoted here as (a), (b), (c), and (d)). These specific configurations are as follows: (1) Component (a), (2) Component (b), (3) Component (b), (4) Components (a) + (b), (5) Components (a) + (c), (6) Components (b) + (c), (7) Components (a) + (b) + (c), (8) Components (a) + (d), (9) Components (b) + (d), (10) Components (a) + (b) + (d). The lighting setup includes: directional light, point light and uniform background lighting, which is introduced to better simulate realistic global illumination. Every scene in PS-Verse is rendered as 20 images, each employing a lighting setup randomly chosen from our ten predefined lighting configurations. An example of such an image set for a single scene is illustrated in Fig. A2.
Figure A6: Our PS-Verse dataset contains ground truth for albedo, metallic, normal, and roughness.
A1.3.4 Comparison with Other Datasets
Here, we primarily compare several training datasets: CyclesPS-Train Ikehata (2018), PS-Wild Ikehata (2022), PS-Mix Ikehata (2023), PS-Uni MS-PS Hardy et al. (2024), and our PS-Verse. For a quantitative comparison of these datasets, please refer to Tab. 1. We now present illustrative qualitative comparisons in Fig. A5. Our comprehensive evaluation, encompassing both qualitative and quantitative aspects, leads us to conclude that PS-Verse is the premier training dataset in terms of quality for the Universal PS task.
A1.3.5 PBR Materials
In addition to ground truth normals, our PS-Verse dataset also provides ground truth maps for albedo, metallic, and roughness, as shown in Fig. A6.
Figure A7: Post-PCA visualization of features extracted by different method encoders for the CowPNG from the DiLiGenT Shi et al. (2018). Metrics displayed to the right of each row is (CSIM/SSIM), higher values indicate higher feature similarity.
A1.4 Additional Discussion
A1.4.1 Why Feature Consistency Matters
@@
Building upon approaches such as UniPS Ikehata (2022) and SDM UniPS Ikehata (2023) that utilize comparable decoder designs, a primary driver for our LINO UniPS is the introduction of an improved encoder. The objective of this encoder is to yield more consistent features, thereby achieving superior results. The direct correlation between greater feature consistency and superior normal reconstruction is not an oversimplification; the decoder’s architecture and capabilities also represent a non-negligible factor in overall performance. Nonetheless, the decoders utilized in UniPS, SDM UniPS and our LINO UniPS are architecturally similar, all adhering to the pixel-sampling paradigm. Consequently, when analyzing the relationship between feature consistency (e.g., as measured by CSIM/SSIM) and normal reconstruction performance in Fig. 1, we group these three methods together to facilitate a more direct comparison of the impact of their respective encoder-derived features.
@@
Ours w/MLP refers to a configuration of our LINO UniPS where the standard decoder is replaced by a simple two-layer MLP. As illustrated in Fig. A7, this variant gives the highest feature similarity. Visual inspection of the PCA plots further reveals that its extracted features have effectively disentangled lighting information. We hypothesize that the superior feature similarity of Ours w/mlp compared to the full LINO UniPS (with its original decoder) stems from the constraints imposed by the weaker MLP decoder; this limited decoder capacity compels the encoder to learn more consistent features to facilitate accurate normal reconstruction. While this enhanced feature consistency from the encoder may not entirely compensate for the reduced representational power of the simpler decoder in terms of absolute normal reconstruction quality (when compared to the full LINO UniPS), Ours w/MLP variant nevertheless significantly outperforms SDM UniPS. This finding strongly corroborates our central hypothesis regarding the critical role of a powerful and well-regularized encoder in achieving effective feature disentanglement and consistency.
@@
Uni MS-PS also demonstrates high feature similarity. However, visual analysis of its features (Fig. A7) suggests that they remain considerably entangled with lighting information. Consequently, we infer that its high reported feature similarity may be more attributable to geometric self-consistency within its representations rather than successful illumination decoupling. Despite this apparent lack of complete feature decoupling, Uni MS-PS often achieves commendable reconstruction results. We attribute this primarily to its multi-scale architecture: beyond the initial stage, each subsequent network stage in Uni MS-PS incorporates predicted normals from the preceding stage as an additional input, effectively leveraging them as a strong geometric prior. This iterative refinement, guided by intermediate normal predictions, places Uni MS-PS in a distinct operational paradigm compared to methods like UniPS, SDM UniPS, and our LINO UniPS.
@@
Furthermore, We need to figure out that Uni MS-PS exhibits certain practical limitations. (a) Its multi-scale nature leads to considerable inference latency, particularly when processing multiple high-resolution input images (e.g., handling 16 images at 4K resolution can extend to about an hour). In contrast, our LINO UniPS method typically completes inference within tens of seconds for similar inputs (see Tab. 6). (b) While Uni MS-PS can reconstruct detailed surface normals, its reliance on potentially lower-resolution training datasets and its patch-based inference mechanism can lead to a loss of global contextual information, sometimes resulting in reconstructions that are locally detailed but globally inconsistent or erroneous (see Fig. 1).
@@
Figure A8: For near-planar objects possessing intricate concave and convex surface details, our LINO UniPS tends to invert the predicted surface normals. The objects are from DiLiGenT-$\Pi$ Wang et al. (2023)
A1.4.2 Limitations
@@
Despite the commendable performance demonstrated by LINO UniPS, certain limitations nonetheless remain, offering avenues for future research.
@@
Secondly, a limitation arises in the estimation of surface normals for near-planar objects that exhibit intricate concave and convex surface details, where the performance of LINO UniPS can be suboptimal. We attribute this primarily to a fundamental challenge inherent in the Universal Photometric Stereo (PS) paradigm: the absence of explicit light source parameters makes it difficult for the network to unambiguously determine the precise illumination direction (e.g., to distinguish between light originating from above versus below the surface). Consequently, this ambiguity can frequently, and in some instances, even lead to the estimated surface normals being inverted (see Fig. A8). In future investigations, we plan to explore whether more sophisticated lighting alignment strategies could effectively mitigate this specific issue.
A1.4.3 The Use of Large Language Models (LLMs)
During the preparation of this manuscript, we utilized Large Language Models (LLMs), including Google’s Gemini, as a writing assistant. The primary application of these models was for language enhancement tasks, such as improving grammar, refining phrasing for clarity, and ensuring stylistic consistency throughout the paper. It is important to note that the core scientific contributions, experimental design, results, and analyses presented herein were conceived and executed solely by the authors. The LLMs’ role was strictly limited to that of a sophisticated tool for polishing the language and presentation of our work.
A1.5 Extended Results
In this section, we present additional results to further demonstrate the capabilities of our LINO UniPS.
A1.5.1 Public Benchmarks
Figure A9: Top row: Example from the input multi-light images. Bottom row: Surface normal map reconstructed by our LINO UniPS.
[…]
Figure A11: Real-world data with masks and corresponding LINO UniPS reconstruction results; data sourced from UniPS Ikehata (2022) and SDM UniPS Ikehata (2023).
A1.5.2 Real Data
Fig. A9 showcases our method’s performance on challenging real-world data. To demonstrate its robustness, the figure includes mask-free examples from two distinct sources: the two leftmost columns show 4K resolution images from the SDM UniPS real data Ikehata (2023), while the rightmost column displays our own $\times960$ captures using a mobile phone.
These results demonstrate that our LINO UniPS is robust to variations in scale and mask-free inputs, while consistently reconstructing fine-grained details.
@@
Fig. A12 presents a comparison of various Universal PS methods on our PS-Verse Testdata. Given that PS-Verse Testdata is a synthetic dataset, ground truth is readily available, facilitating precise quantitative evaluation. The results clearly demonstrate that our LINO UniPS significantly outperforms contemporary approaches, including Uni MS-PS Hardy et al. (2024), SDM UniPS Ikehata (2023), and UniPS Ikehata (2022).